Cauchy. 4. (4 points) Suppose X ~ (a) Find P(X E [0, 8]). (Hint: tan-(a) = 1Haz.) (b) Find P(3X + 5 € [0, 8]). 1+a²
Answer #1

8: Parameter of x is not geven, So assceme x as Standard Cauchy distribution i.e. Xin Cauchy (6,9) f(x) = 2 - F(x) = h / 2 +